OpenStudy (anonymous):

Jade hiked 1 1/4 mile on the trail. Gretta hiked 1 1/2 times as far on the trail as Jade. How far did Gretta hike on the trail?

OpenStudy (anonymous):

2 3/4

OpenStudy (anonymous):

because 1/2 is 2 1/4 so 1 1/4 plus 2/4 and 1 equals 2 3/4

OpenStudy (anonymous):

you sure ?

OpenStudy (texaschic101):

its multiplication..not addition

OpenStudy (texaschic101):

do you know how to turn mixed fractions into improper fractions ?

OpenStudy (anonymous):

put the bottom number on top

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

no....let me show you. I will do one for you..then you can do the other one. 1 1/4 step 1 : take the whole number and multiply it by the denominator 1 * 4 = 4 step 2 : add that number to the numerator 4 + 1 = 5 step 3 : now put that number over your original denominator 5/4 so 1 1/4 = 5/4 can you do 1 1/2 ?

OpenStudy (anonymous):

@texaschic101 im from texas to

OpenStudy (texaschic101):

cool.....now can you turn 1 1/2 into an improper fraction...following the steps I put above

OpenStudy (texaschic101):

hello ??

OpenStudy (texaschic101):

ok...well...tag me (if you want) when you are ready to proceed